eeR papers as stableah No Question to instructor! If you belive t statement; use your assumption to solve the problem. Breaking the e Prob-1 Probability that cars have been assembled in a particular plant as follows: Pl plant I )-20% , P(plam2)-24% , P(plant3)-25% , P(plant4)-31 % Total number of cars manufactuted by all plants: 1.2 million/year Number of cars of no warranty claims manufactured by plantl: 223200/year ty claim rate of the cars assembeled in plantl? Ans: Prob-2 The diameter of metal cylinder has a probability density function of fx)-1.5-6(x-50. 1000 metal cylinders delivered to engine assembely plant... How many cylinders' dia Ans: Prob-3 A company makes electric motors. The probability an electric motor is defective is 5% What is the probability that a sample of 100 electric motors will contain no defective m Ans Sol Prob-4 Pressurized air are being used in a manufacturing company in its daily operation. There a compressors, (S1 ,S2) and one-reciprocal type compressor,(R). Each compressor, at any g is denoted to reciprocal compressor (R) that is always off. Event B is denoted that at least that all outcomes in event B are equaly likely. If P(AnB)-45% and P(AUB)-93% Comp Ans Sol:
DIODIOn!. BPeaking the exam pölicy will be treated as cheat Prob-1 Probability that cars have been assembled in a particular plant as follows: P(plant 1)-20% , P(plant2)-24% , P(plam3)-25%. P(plant4)-31% Total number of cars manufactuted by all plants: 1,2 million year Number of cars of no warranty claims manufactured by planti: 223200/year eled in plantl? Ans Prob-2 The diameter of metal eylinder has a probability density function of fix-1.5-6ix-50.0P [mm] 1000 mctal cylinders delivered to engine assembely plant... How many cylinders' diameters, within4 .8s Ds 50.3 mm? Ans: Prob-3 A company makes electric motors. The probability an electric motor is defective is 5% What is the probability that a sample of 100 electric motors will contain no defective motor? Use just one decimal point! Ans Sol Prob-4 Pressurized air are being used in a manufacturing company in its daily operation. There are three compressors in service; two-scre compressors, (S1 ,S2) and one-reciprocal type compressor (R). Each compressor, at any given time, is either off (O)%, or on (1). Ass is denoted to reciprocal compressor (R) that is always off. Event B is denoted that at least one of the screw type of compressor is a that all outcomes in event B are equaly likely. If P(AnB 45% and P(AUB)-93%. Comput ieprobabi itthat ILcompressorsr Ans Sol:
Answer #1

1)
number of car assemble in plant 1 = 0.20 * 1.2*10^6
= 240000

hence warranty claim rate = 1 - 223200/240000
= 0.07

2)

P(49.8 < X < 50.3) =

50.3 1.5-6 (x-50)2 dx = 0.68 49.8

hence we expect 1000 *0.68 = 680 cylinders with diameter between 49.8 and 50.3

3)

probability that there are 0 defective

= (1-p)^n

= 0.95^100

=

0.00592052922

Please rate

Please post one question at a time